# No Train but Gain: Language Arithmetic for training-free Language Adapters enhancement

### v1

@@ 4.3 Language Arithmetic evaluation
This evaluation assumes that the target language adapter exists in the adapter pool. We test two cases, i.e. $LA(T,en)$ and $LA(T,REL)$, where $REL$ is once again a related language and $T$ is the target language. The results are presented in Table 2. Compared to the baseline direct application of a target language adapter (i.e. MAD-X method), the gains are not as significant as in the case of the zero-shot scenario. Given that a target language adapter is trained on a significant corpus, it gives less room for improvement (this is not the case in the low-resource regimes, as shown in the following Section). However, considering the cost-to-performance ratio and the minimal fatigue that our post-processing method enforces on existing MAD-X pipelines, we can see a constant but minor gain.
Notably, during the evaluation of XLM-R NLI, we observed that the outcomes remain consistent across different methods (e.g. $67.8$ for Swahili). This behaviour is a built-in property of the proposed method. The method has a fail-safe mechanism, which keeps the extreme values during the evaluation of $\lambda\in[0,1]$. Suppose the method does not find a better lambda different than the extreme values (i.e. $\lambda\in\{0,1\}$). In that case, it means that we are using a proxy language directly, and the language arithmetic did not provide an improvement (but kept the final result the same).
Table 2: Results over all tested languages of the improvement of a target language adapter (denoted as $T$, i.e. the MAD-X method [29]) via language arithmetic. We use both English ($en$) and a related language ($REL$) as a right side of a $LA$ function. Additionally, we report a mix ($MIX$) results where we choose better arithmetic combination for each language.
$T\rightarrow$ ar bg de el es fr hi ru sw tr ur vi zh AVG
NER (F1 score)
mBERT $T$ 37.03 79.30 78.50 74.79 74.83 78.32 68.50 65.95 73.65 71.70 40.68 72.64 36.77 65.59
$LA(T,en)$ 37.03 80.11 78.50 75.56 74.64 78.24 66.13 65.95 74.72 72.37 40.68 73.12 39.62 65.90
$LA(T,REL)$ 37.03 79.64 78.50 75.12 74.83 78.26 68.09 66.04 74.50 71.89 40.68 72.88 39.21 65.90
$LA(T,MIX)$ 37.03 80.11 78.50 75.56 74.83 78.26 68.09 66.04 74.72 72.37 40.68 73.12 39.62 66.07
[…]
$LA(T,en)$ 51.32 79.60 77.56 78.65 78.86 76.90 72.27 62.37 74.84 76.28 63.96 66.68 27.55 68.22
$LA(T,REL)$ 51.23 80.16 77.59 78.50 79.06 76.70 71.72 62.34 74.53 76.28 63.96 66.18 28.10 68.18
$LA(T,MIX)$ 51.32 80.16 77.59 78.65 79.06 76.90 72.27 62.37 74.84 76.28 63.96 66.68 28.10 68.32
NLI (Accuracy)
mBERT $T$ 63.41 67.35 67.88 65.23 71.72 72.08 61.24 66.01 58.98 62.99 58.54 68.16 67.64 65.48
$LA(T,en)$ 63.41 67.35 68.46 65.23 72.24 72.24 61.24 66.01 59.16 62.81 58.10 68.54 68.22 65.62
$LA(T,REL)$ 63.41 67.35 68.80 65.09 71.12 72.28 61.08 65.65 58.96 62.99 58.48 68.94 67.90 65.54
$LA(T,MIX)$ 63.41 67.35 68.80 65.23 72.24 72.28 61.24 66.01 59.16 62.99 58.48 68.94 68.22 65.72
[…]
$LA(T,en)$ 70.36 76.59 75.37 74.33 77.41 77.03 69.34 74.73 67.80 71.54 66.03 74.19 73.47 72.94
$LA(T,REL)$ 71.32 76.59 75.35 74.97 77.13 76.97 68.84 74.73 67.80 71.12 66.15 73.85 72.71 72.89
$LA(T,MIX)$ 71.32 76.59 75.37 74.97 77.41 77.03 69.34 74.73 67.80 71.54 66.15 74.19 73.47 73.07
QA (F1 score)
mBERT $T$ 54.46 - 71.17 51.14 73.45 - 48.80 66.92 - 49.27 - 65.85 54.10 59.46
$LA(T,en)$ 55.84 - 71.13 53.41 72.83 - 50.36 67.62 - 48.85 - 63.54 54.15 59.75
$LA(T,REL)$ 55.42 - 70.99 54.71 73.16 - 48.80 67.07 - 49.31 - 64.94 54.94 59.93
$LA(T,MIX)$ 55.84 - 71.13 54.71 73.16 - 50.36 67.62 - 49.31 - 64.94 54.94 60.22
[…]
$LA(T,en)$ 67.61 - 72.71 73.44 75.60 - 67.96 72.02 - 66.57 - 74.07 63.24 70.36
$LA(T,REL)$ 68.00 - 71.62 73.66 75.71 - 67.96 72.03 - 67.01 - 74.12 64.85 70.55
$LA(T,MIX)$ 68.00 - 72.71 73.66 75.71 - 67.96 72.03 - 67.01 - 74.12 64.85 70.67
@@ 4.4 Low resource evaluation
Figure 2: NER (left) and NLI (right) evaluation of a set of adapters trained on a Wikipedia subset showcases that language arithmetic $LA(T,en)$ (green, dotted line) provides significant gains when compared against direct usage of the adapter (violet, solid line), especially in a very low-resource regime. The x-axis represents the token budget of each trained language adapter.
Training a language adapter might be troublesome for high-resourced languages due to massive corpora requiring significant computational resources. [^6] Although in our experimental setup, we train each adapter for the same number of steps and choose the best checkpoint based on the validation performance, for low-resourced languages, one could apply an early stopping mechanism in a production-level pipeline. On the other hand, in most languages, we lack data to train a strong language adapter, i.e. language-specific corpora might be either too small or unavailable [15]. We investigate whether language arithmetic can help in such cases. We test our solution in three cases and define three (actual and simulated) evaluation scenarios:
  • Assamese (as) - low resource language, additionally not used in the pre-training of mBERT and XLM-R,
@@
We train a series of language adapters with different token budgets for each language, from 10k to 10M (or 1B for French; we limit this particular study to the XLM-R model). Afterwards, we compare the usage of such adapters directly against language arithmetic with three adapters (we use $LA(T,en)$, where $T\in\{as,sw,fr\}$).
Figure 2 presents the results of the evaluation performed on the downstream tasks. The most gain is visible in the most challenging scenario, during the evaluation on the Assamese dataset. In this case, the backbone MLLMs did not encounter the language during the pre-training phase. Although the difference becomes less pronounced in the NER task as we approach the limits of available data, there remains a significant margin for NLI - the difference can be explained by the overlap in the corpora (Wikipedia) between NER and language adapter training tasks, following the findings of Gururangan et al. [13]. For Swahili, where the language is part of the pre-training, the flattening effect begins earlier and affects both tasks. Nevertheless, leveraging language arithmetic still yields improvements.
The simulated case of French showcases that even with a relatively weak language adapter (trained on 10k tokens), the language arithmetic can restore existing knowledge and results in high performance for the language. Moreover, comparing the adapters trained with a different token budget, the results remain similar, without significant fluctuations. We believe that this phenomenon happens because the Multilingual Pre-trained Language Model has seen a much higher amount of French in the pre-training procedure than Swahili (over $35$ times more tokens in XLM-R pre-training; moreover, French is in the top $15$ represented languages). Therefore, even undertrained French adapters have a relatively easy task once they are merged with a robust English adapter. In practical terms, this finding allows us to prototype new languages quicker by estimating the possible end product quality or might serve as an intermediate solution (until the full-corpora adapter is trained).
@@ 5.1 Lambda impact
Our study estimates the $\lambda$ parameter with a small step ($0.05$). Depending on multiple variables that include, among others, model and evaluation dataset sizes or number of languages, running $20$ evaluations might be troublesome (especially when using neural-based metrics, e.g. COMET [38]). Therefore, we analysed the potential impact of choosing a suboptimal lambda with a decrease in evaluation count. The breakdown includes a subset of languages on both tasks (using the XLM-R version). We chose the zero-shot scenario where we performed language arithmetic between English and related language adapters.
In Figure 3, we plot the validation scores with the corresponding baselines, that is, the scores of using directly the adapters. The dotted lines are based on $\lambda=0$ or $\lambda=1$ for clarity, meaning we exclusively use the arithmetic equation’s left or right side (i.e., a specific language). In most cases, a subset of valid $\lambda$ values would improve over the baselines. Moreover, the analysis reveals that a coarser evaluation (e.g., with a step of $0.1$ or $0.2$) would be sufficient, reducing the required number of performed tests up to four times while maintaining most of the improvement. At last, setting the default $\lambda=0.5$ would be near optimum for the analysed subset.
Figure 3: Interpolation of $\lambda$ values for the zero-shot XLM-R scenario (NER - top, NLI - bottom; for QA see Appendix A) on the validation dataset. The horizontal dashed lines represent the baseline scores for both languages used in language arithmetic.
@@ 5.2 Language Relatedness
The results are presented in Table 3 and show that given a related language (in this case, defined as coming from the same language family), there might be very little difference in which one we will use. The relative difference between the best and the worst language reaches only a $0.40$ F1 score in the NER task and the same value in terms of accuracy points in NLI. Nevertheless, the more sophisticated and hand-crafted language choice would improve the results from Table 2.
Table 3: Impact of language relatedness on the language arithmetic. We compare different Romance languages as a right side of $LA$ equation, i.e. $l2$ (both tasks use XLM-R model).
$LA(l_{1}\downarrow,l_{2}\rightarrow)$ ca es fr it pr ro
Eval language NER
es 78.77 - 78.96 79.06 78.91 78.66
fr 76.63 76.92 - 76.70 76.80 76.80
@@
es 77.07 - 77.23 77.13 77.07 76.83
fr 76.77 76.99 - 76.97 76.95 77.07
5.3 Language vs task vectors
Task vectors exhibit high sparsity and orthogonality, as Ilharco et al. [19] observed. While the former characteristic can be denoted in language vectors (Figure 4), the latter displays different properties, in contrast to task vectors. In Figure 5, we visualise the cosine similarity between evaluated language vectors of language adapters. Notably, the minimal cosine similarity ($0.2$) surpasses the maximum ($0.18$) reported by previous research in the task space [19]. Interestingly, most pairs in the task space oscillate within the range of $0.01$ to $0.03$. At the same time, language vectors surpass $0.2$ in each case, indicating that the orthogonality aspect is an inherent property of task adapters.
@@ 5.3 Language vs task vectors
Based on the cosine similarity observation, we investigated one of the recent task arithmetic extensions, Ties-Merging [44]. This work introduces a three-step algorithm that prevents different parameter interferences, improving upon task arithmetic. The algorithm decreases the cosine similarity via a pruning step and alignment of parameter signs to perform arithmetic only on relevant parameters to the merged tasks. On the experimental details note, as Ties-Merging operates on averaging, not addition, we utilise a different lambda range during validation (as suggested by Yadav et al. [44], $\lambda\in[0.8,1.8]$), and we set Top-Kto the default value of $20$ and additionally to $80$.
@@
We report the comparison in the zero-shot setting on the NER task (XLM-R version) in Table 4. The Ties-Merging decreases the results significantly; it is not only outperformed by language arithmetic in every language pair but also by other baselines (see Table 1). Moreover, we note that the pruning operation has the reverse effect; higher pruning (i.e. keeping Top-Klower) decreases the performance (in contrast to task vectors) by making language vectors more sparse and, hence, closer to orthogonal.
Table 4: Ties-Merging evaluation in the zero-shot setup on the NER task (XLM-R version). In the case of language arithmetic, where the language vectors have a higher overlap (i.e. higher cosine similarity), removing parameter interference decreases the overall performance.
@@
LA 61.20
Ties-Merging (Top-K20) 56.40
Ties-Merging (Top-K80) 59.88
@@
6 Related Work
Knowledge composition from multiple, independently trained adapters has been widely discussed in the literature. However, unlike our work, the solutions require substantial changes to the vanilla adapter setup. The previous work either requires additional parameters to learn a parameterised composition function/a gating module to combine/steer the flow through the suitable adapter(s), or needs a specific training procedure that increases the complexity of the overall solution or, in most cases, both [33, 30, 22, 26, 9, 20, 41]. Moreover, to prevent specifically negative interference, hyper-adapters [7] were proposed using hyper-networks [14], and Ansell et al. [3] applied sparse fine-tuning to compose task and language masks. Unlike the prior studies mentioned earlier, our attempt is training-free and does not modify the base architecture. The most conceptually similar work is proposed by Chronopoulou et al. [8]; however, they operate on the notion of sample similarity to a subset of domains in a domain adaptation regime. Additionally, concurrent to our work, Parović et al. [27] show intial potential of task arithemtic in cross-lingual transfer based on a full fine-tuning setup. At last, we denote the rise of task arithmetic use cases, e.g. vision tasks or cross-task generalisation [39, 18].
7 Conclusion
We have proposed language arithmetic, which enhances language adapters based on task arithmetic learning via addition. It is a training-free method and functions as a post-processing technique for the adapters. Our experiments have shown that LA is particularly beneficial in zero-shot and low-resourced cases. At last, we highlight the differences between language and task arithmetic.
In our future work, we plan to extend language arithmetic by incorporating additional components into the sum. Additionally, we aim to adapt other elements of the task arithmetic framework, initially designed for tasks, i.e. task analogies and forgetting via negation, to a multilingual setup with an analysis of the differences between multi-task and multilingual arithmetic context. Furthermore, we will evaluate LA’s performance on various non-classification tasks.
8 Limitations
Our work was tested on English-centric task training and could be extended to different languages with more PEFT methods. Moreover, applying multi-source training based on the work of Ansell et al. [4] could provide better robustness of the task adapters and a more thorough analysis. At last, our analysis of the differences between task and language vectors is preliminary, without definitive conclusions.
Appendix A Lambda impact - QA
Figure 6 presents the analysis of lambda impact for the Question Answering task. For details, refer to Section 5.1.
Figure 6: Interpolation of $\lambda$ values for the zero-shot QA XLM-R scenario on the validation dataset. The horizontal dashed lines represent the baseline scores for both languages used in language arithmetic.
Appendix B Related languages choice
We derived related languages for the language adapters based on the available pool, i.e. we reused the adapters from other languages from the task (we trained two bonus ones, Persian and Italian, apart from the analysis in Section 5.2). While the related language could have been more suitable in a few cases, there is always a trade-off between relatedness, criterion for relatedness and available corpora. Moreover, the perfectly related language matching would be an unfair comparison as it would assume the availability of the ideal adapter’s pool, which is unlikely in practice. Table 5 presents the related language choice used in our experiments.
Table 5: Languages used in the experiments with corresponding related languages. In the most cases, we used a related language from the already existing pool (i.e. evaluated languages, the first rows labelled “Lang.”); however, we trained some additional ones, including Italian and Persian.
Lang. ar bg de el es fr hi
Related Lang. fa ru fr de it it ur
[…]
Related Lang. bg ar de hi zh vi

### v2

@@ 4.3 Improving existing language adapters
The drop in performance of language arithmetic compared to the zero-shot use case is not surprising. Given that a target language adapter is trained on a significant corpus, it gives less room for improvement (this is not the case in the low-resource regimes, as shown in the following Section). This is also a potential explanation of a different behaviour between the evaluated MLLMs, considering the overall more robust performance of XLM-R over mBERT. However, considering the cost-to-performance ratio and the minimal fatigue that our post-processing method enforces on existing MAD-X pipelines, we can see a constant gain on average across all the experiments and training runs.
Figure 3: Variants of language arithmetic compared to the MAD-X method in the use-case to improve an existing target language adapter. The values above bars present a difference between a better LA setup and the MAD-X framework for the XLM-R model (see Figure 7 for mBERT).
@@ 4.4 Low resource evaluation
((a)) NER
((b)) NLI
Figure 4: NER and NLI evaluation of a set of adapters trained on a Wikipedia subset showcases that language arithmetic $LA_{t}(t,en)$ (green, dotted line) provides significant gains when compared against direct usage of the adapter (violet, solid line), especially in a very low-resource regime. The x-axis represents the token budget of each trained language adapter.
Training a language adapter might be troublesome for high-resourced languages due to massive corpora requiring significant computational resources. [^6] Although in our experimental setup, we train each adapter for the same number of steps and choose the best checkpoint based on the validation performance, for low-resourced languages, one could apply an early stopping mechanism in a production-level pipeline. On the other hand, in most languages, we lack data to train a strong language adapter, i.e. language-specific corpora might be either too small or unavailable Haddow et al. (2022). We investigate whether LA can help in such cases. We test our solution in three cases and define three (actual and simulated) evaluation scenarios:
  • Assamese (as) - low resource language, additionally not used in the pre-training of a base MLLM,
@@
We train a series of language adapters with different token budgets for each language, from 10k to 10M (or 1B for French; we limit this particular study to the XLM-R model). Afterwards, we compare the usage of such adapters directly against language arithmetic with three adapters (we use $LA_{t}(t,en)$, where $t\in\{as,sw,fr\}$).
Figure 4 presents the results of the evaluation performed on the downstream tasks. The most gain is visible in the most challenging scenario, during the evaluation on the Assamese dataset. In this case, the backbone MLLMs did not encounter the language during the pre-training phase. Although the difference becomes less pronounced in the NER task as we approach the limits of available data, there remains a significant margin for NLI - the difference can be explained by the overlap in the corpora (Wikipedia) between NER and language adapter training tasks, following the findings of Gururangan et al. (2020). For Swahili, where the language is part of the pre-training, the flattening effect begins earlier and affects both tasks. Nevertheless, leveraging language arithmetic still yields improvements.
The simulated case of French showcases that even with a relatively weak language adapter (trained on 10k tokens), the language arithmetic can restore existing knowledge and results in high performance for the language. Moreover, comparing the adapters trained with a different token budget, the results remain similar, without significant fluctuations. We believe that this phenomenon happens because the MLLM has seen a much higher amount of French in the pre-training procedure than Swahili (over $35$ times more tokens in XLM-R pre-training; moreover, French is in the top $15$ represented languages). Therefore, even undertrained French adapters have a relatively easy task once they are merged with a robust English adapter. In practical terms, this finding allows us to prototype new languages quicker by estimating the possible end product quality or might serve as an intermediate solution (until the full-corpora adapter is trained).
@@ 5.1 Lambda impact
Our study estimates the $\lambda$ parameter with a small step ($0.05$). This analysis investigates how sensitive this parameter is in the language arithmetic. Depending on multiple variables that include model and evaluation dataset sizes or a number of languages, running $20$ evaluations might be costly (especially when using neural-based metrics, e.g. COMET Rei et al. (2020)). Therefore, we analysed the potential impact of choosing a suboptimal lambda with a decrease in evaluation count. The breakdown includes a subset of languages on both tasks (using the XLM-R as a base model). We chose the zero-shot scenario where we performed LA between English and related language adapters.
In Figures 5 and 8, we plot the validation scores with the corresponding baselines, that is, the scores of using directly the adapters. The dotted lines are based on $\lambda=0$ or $\lambda=1$ for clarity, meaning we exclusively use the arithmetic equation’s left or right side (i.e., a specific language). In most cases, a subset of valid $\lambda$ values would improve over the baselines. Moreover, the analysis reveals that a coarser evaluation (e.g., with a step of $0.1$ or $0.2$) would be sufficient, reducing the required number of performed tests up to four times while maintaining most of the improvement. At last, setting the default $\lambda=0.5$ would be near optimum for the analysed subset.
Figure 5: Interpolation of $\lambda$ values for the zero-shot XLM-R scenario (NER, for NLI and QA see Appendix D) on the validation dataset. The horizontal dashed lines represent the baseline scores for both languages used in language arithmetic.
@@ 5.2 Language relatedness
$LA_{l_{1}}(l_{1}\downarrow,l_{2}\rightarrow)$ ca es fr it pt ro
Eval language NER
es 73.82 - 75.06 74.61 74.90 73.68
fr 75.74 75.76 - 75.82 75.59 75,79
[…]
es 78.23 - 78.04 77.96 77.94 78,02
fr 77.95 77,65 - 77.70 77,47 77,60
Table 1: Impact of language relatedness on the language arithmetic. We compare different Romance languages as a right side of $LA$ equation, i.e. $l2$ (both tasks use XLM-R model). We report an average over three runs.
@@
The results are presented in Table 1 and show that given a different related language (in this case, defined as coming from the same language family), there are minor scores fluctuation. The relative difference between the best and the worst language reaches around $1$ F1 score in the NER task and around $0.3$ in terms of accuracy points for the NLI task. This experiment indicates that a more sophisticated or hand-crafted language choice would improve the downstream results presented in Figure 5. However, it also shows that there is no free lunch, and results depend on a downstream task. For example, for Spanish evaluation, a French adapter is trained on the largest out of the listed languages raw corpora; therefore, for the NER task, it can leverage a bigger pool of seen during training Named Entities (at times language-independent or similar across languages) and perform the best given more data, even when there are closer related languages (according to methodology from Section 4.1.1 and Indo-European languages family tree).
6 Related Work
Knowledge composition from multiple, independently trained adapters has been widely discussed in the literature. However, unlike our work, the solutions require substantial changes to the vanilla adapter setup. The previous work either requires additional parameters to learn a parameterised composition function/a gating module to combine/steer the flow through the suitable adapter(s), or needs a specific training procedure that increases the complexity of the overall solution or, in most cases, both Pham et al. (2020); Pfeiffer et al. (2021); Lee et al. (2022); Parović et al. (2022); Chronopoulou et al. (2023b); Klimaszewski et al. (2023); Wang et al. (2023). Moreover, to prevent specifically negative interference, hyper-adapters Baziotis et al. (2022) were proposed using hyper-networks Ha et al. (2017), and Ansell et al. (2022) applied sparse fine-tuning to compose task and language masks. Unlike the prior studies mentioned earlier, our attempt is training-free and does not modify the base architecture. The most conceptually similar work is proposed by Chronopoulou et al. (2023a); however, they operate on the notion of sample similarity to a subset of domains in a domain adaptation regime. Additionally, concurrent to our work, Parović et al. (2024) show intial potential of task arithemtic in cross-lingual transfer based on a full fine-tuning setup. However, in our work we focus on PEFT methods with additional, in-depth analysis. At last, we denote the rise of task arithmetic use cases, e.g. vision tasks or cross-task generalisation Stoica et al. (2023); Huang et al. (2024).
7 Conclusion
We have proposed language arithmetic, which enhances language adapters based on task arithmetic learning via addition. It is a training-free method and functions as a post-processing technique for MAD-X. Our experiments have shown that LA is particularly beneficial in a zero-shot scenario, where we do not have access to a target language adapter. At last, we highlight the differences between language and task arithmetic.
In our future work, we plan to extend language arithmetic by incorporating more components into the sum. Additionally, we aim to adapt other elements of the task arithmetic framework, i.e. task analogies and forgetting via negation, to a multilingual setup with an analysis of the differences between multi-task and multilingual arithmetic context. Furthermore, we will evaluate LA’s performance on various non-classification tasks.
8 Limitations
Our work was tested on English-centric task training and could be extended to different languages with more PEFT methods. Moreover, applying multi-source training based on the work of Ansell et al. (2023) could provide better robustness of the task adapters and a more thorough analysis.
Appendix A Related languages
We present the list of related languages used in our experiments in Table 2 (details in Section 4.1.1).
Lang. ar bg de el es fr hi
Related sw ru fr es fr es ur
@@ Appendix A Related languages
Related bg ar bg hi ru ar
Table 2: Languages used in the experiments with corresponding related languages. Details are provided in Section 4.1.1.
Appendix B Zero-shot evaluation
Figure 6: Zero-shot mBERT language arithmetic evaluation, where one side of the arithmetic is an English adapter, and the other is related to the target language adapter (e.g. French for Spanish - $LA_{es}(en,fr)$). The values above bars present a relative difference to a better proxy.
Figure 6 presents the results of the experiments described in Section 4.2 for the mBERT model.
Appendix C Improving existing language adapters
Figure 7 presents the results of the experiments described in Section 4.3 for the mBERT model.
Figure 7: Variants of language arithmetic compared to the MAD-X method in the use-case to improve an existing target language adapter. The values above bars present a difference between a better LA setup and the MAD-X framework for the mBERT model.
Appendix D Lambda impact - NLI and QA
Figure 8 presents the analysis of lambda impact for NLI and QA tasks. For details, refer to Section 5.1.
Figure 8: Interpolation of $\lambda$ values for the zero-shot NLI and QA XLM-R scenario on the validation dataset. The horizontal dashed lines represent the baseline scores for both languages used in language arithmetic.
Appendix E Language vs task vectors
Task vectors exhibit high sparsity and orthogonality, as Ilharco et al. (2023) observed. While the former characteristic can be denoted in language vectors (Figure 9), the latter displays different properties, in contrast to task vectors. In Figure 10, we visualise the cosine similarity between evaluated language vectors of language adapters. Notably, the minimal cosine similarity ($0.19$) surpasses the maximum ($0.18$) reported by previous research in the task space Ilharco et al. (2023). Interestingly, most pairs in the task space oscillate within the range of $0.01$ to $0.03$. At the same time, language vectors surpass $0.2$ in almost each case, indicating that the orthogonality aspect is an inherent property of task adapters.
@@ Appendix E Language vs task vectors
Based on the cosine similarity observation, we investigated one of the recent task arithmetic extensions, Ties-Merging Yadav et al. (2023). This work introduces a three-step algorithm that prevents different parameter interferences, improving upon task arithmetic. The algorithm decreases the cosine similarity via a pruning step and alignment of parameter signs to perform arithmetic only on relevant parameters to the merged tasks. On the experimental details note, as Ties-Merging operates on averaging, not addition, we utilise a different lambda range during validation (as suggested by Yadav et al. (2023), $\lambda\in[0.8,1.8]$), and we set Top-Kto the default value of $20$ and additionally to $80$.
@@
We report the comparison in the zero-shot setting on the NER task (XLM-R version) in Table 3. The Ties-Merging decreases the results significantly compared to the default language arithmetic. Moreover, we note that the pruning operation has the reverse effect; higher pruning (i.e. keeping Top-Klower) decreases the performance (in contrast to task vectors) by making language vectors more sparse and, hence, closer to orthogonal.
@@
LA 60.54
Ties-Merging (Top-K20) 52.94
Ties-Merging (Top-K80) 57.57
Table 3: Ties-Merging evaluation in the zero-shot setup on the NER task (XLM-R version, averaged over three runs and all evaluated languages). In the case of language arithmetic, where the language vectors have a higher overlap (i.e. higher cosine similarity), removing parameter interference decreases the overall performance.